# Characteristics of Political Misinformation Over the Past Decade

### v1

@@ 4 Misinformation Source Modality Trends
As a result of the multimodal nature of contemporary misinformation, researchers should no longer rely exclusively on algorithms associated with text (natural language processing) or images/video (computer vision) alone, and transition to multimodality versions available models. Alternatively, they may choose to deploy independent models to detect and mitigate for each modality, but this may require additional complexity.
@@ 6 Misinformation Topics Across Years
2021 COVID (359), Election Fraud (94), Joe Biden (58), Hillary Clinton (55), Capital Riot (54) 17
@@
Exploring topics within each year, it is apparent that some topics are heavily aligned with current events and themes that are temporally localized (e.g., Capital Riot). However, there are some topics that recur across years and the next section will explore those topics further.
@@ 8 Conclusions
This paper also uncovered several recurring topics of misinformation (Figure 5) spanning seven categories (Table 4), which provides insight into temporally invariant themes of misinformation. Taken in the context that misinformation is emotionally charged, these categories could suggest that people tend to share inaccurate statements around information they fear or don’t understand (Science and Medicine, Crime, Religion), impacts them directly (Policy, Election Integrity, Economic) or Public Figures who are salient in their daily lives.

### v2

@@ 4 Misinformation Source Modality Trends
As a result of the multimodal nature of contemporary misinformation, researchers should no longer rely exclusively on algorithms associated with text (natural language processing) or images/video (computer vision), and transition to multimodality versions available models. Alternatively, they may choose to deploy independent models to detect and mitigate for each modality, but this may require additional complexity.
@@ 6 Misinformation Topics Across Years
2021 COVID (359), Election Fraud (94), Joe Biden (58), Hillary Clinton (55), Capitol Riot (54) 17
@@
Exploring topics within each year, it is apparent that some topics are heavily aligned with current events and themes that are temporally localized (e.g., Capitol Riot). However, there are some topics that recur across years and the next section will explore those topics further.
@@ 8 Conclusions
This paper also uncovered several recurring topics of misinformation (Figure 5) spanning seven categories (Table 4), which provides insight into temporally invariant themes of misinformation. Taken in the context that misinformation is emotionally charged, these categories could suggest that people tend to share inaccurate statements around information they fear or don’t understand (Science and Medicine, Crime, Religion), that impacts them directly (Policy, Election Integrity, Economic) or Public Figures who are salient in their daily lives.